# Addressing a fundamental limitation in deep vision models: lack of spatial attention

### v2

@@ Abstract
The primary aim of this manuscript is to underscore a significant limitation in current deep learning models, particularly vision models. Unlike human vision, which efficiently selects only the essential visual areas for further processing, leading to high speed and low energy consumption, deep vision models process the entire image. In this work, we examine this issue from a broader perspective and propose a solution that could pave the way for the next generation of more efficient vision models. Basically, convolution and pooling operations are selectively applied to altered regions, with a change map sent to subsequent layers. This map indicates which computations need to be repeated. The code is available at https://github.com/aliborji/spatial_attention.
@@ 1 Motivation
In the proposed approach, computation is performed on demand. One advantage of this method is that it can be applied solely during inference. The model can be trained using a GPU and then optimized using this approach to enhance inference efficiency.
@@ 2 Related work
3 A potential solution
In the proposed approach, convolution and pooling operations are selectively applied to altered regions, with a change map sent to subsequent layers. This map informs those layers about which computations need to be repeated. Each layer communicates changes so the next layer knows what it needs to recompute, and this process continues until the final layer. To achieve this, each layer must remember its last computation to avoid redundant processing.
@@ 4 Experiments and results
4.1 Experiment I: Processing repeated versions of the same image
@@
4.2 Experiment II: Processing shifted versions of the same image
This experiment is similar to Experiment I, except each of the 11 digits is shifted rightward by one pixel at a time, resulting in 110 images in total. This method causes some regions of the image to remain the same while others change. As a result, CNN3 is slower here compared to its speed in Experiment I because it needs to recompute more information due to the increased amount of changed content.
@@ 4.2 Experiment II: Processing shifted versions of the same image
also mention if we have a tensor of size W x H x C channels (say C=100), and if your kernel size is KxK, then you save KxkxC operations are saved.
we were mainly concerned with inference time but maybe this can be also applied for training. finally applying this idea for other aritechtures and tasks such as face detecrtion (laptop camera etc) can be interesting.
@@ 5 Conclusion
We highlighted a key issue with existing deep learning approaches and proposed a simple solution that can be integrated into current models or used to design new models with inherent attention and memory mechanisms. This work serves as a proof of concept and can be applied to other problems such as object detection, scene segmentation, and action recognition. It also has the potential to help address adversarial examples [2]. This method is particularly effective when the input has higher resolution.
@@ 6 Appendix
Convolution and pooling modules of CNNs are shown in Figures 78, and  9.

### v3

@@ Abstract
The primary aim of this manuscript is to underscore a significant limitation in current deep learning models, particularly vision models. Unlike human vision, which efficiently selects only the essential visual areas for further processing, leading to high speed and low energy consumption, deep vision models process the entire image. In this work, we examine this issue from a broader perspective and propose two solutions that could pave the way for the next generation of more efficient vision models. In the first solution, convolution and pooling operations are selectively applied to altered regions, with a change map sent to subsequent layers. This map indicates which computations need to be repeated. In the second solution, only the modified regions are processed by a semantic segmentation model, and the resulting segments are inserted into the corresponding areas of the previous output map. The code is available at https://github.com/aliborji/spatial_attention.
@@ 1 Motivation
In the first proposed approach, computation is performed on demand. One advantage of this method is that it can be applied solely during inference. The model can be trained using a GPU and then optimized using this approach to enhance inference efficiency. In the second approach, only the modified regions are processed by a semantic segmentation model, and the resulting segments are inserted into the corresponding areas of the previous output map.
@@ 2 Related work
3 Potential solution I
In this approach, convolution and pooling operations are selectively applied to altered regions, with a change map sent to subsequent layers. This map informs those layers about which computations need to be repeated. Each layer communicates changes so the next layer knows what it needs to recompute, and this process continues until the final layer. To achieve this, each layer must remember its last computation to avoid redundant processing.
@@ 3.1 Experiments and results
Experiment I: Processing repeated versions of the same image
@@
Experiment II: Processing shifted versions of the same image This experiment is similar to Experiment I, except each of the 11 digits is shifted rightward by one pixel at a time, resulting in 110 images in total. This method causes some regions of the image to remain the same while others change. As a result, CNN3 is slower here compared to its speed in Experiment I because it needs to recompute more information due to the increased amount of changed content.
@@
4 Potential solution II
The second method involves semantic segmentation, often regarded as a supertask. This is because the capability to label each individual pixel can generalize to other tasks, such as object recognition and object detection. (Instance segmentation, however, is even more comprehensive since it distinguishes between instances of the same object.) The idea here is to feed only the modified regions into the segmentation model. A benefit of some models, especially those based on U-Net architectures [9], is that they can process inputs of any resolution. After obtaining the output, it can be accurately positioned in the prior output map.
Figure 7 provides a more detailed illustration. In the middle panel, the regions with changes from the previous frame (e.g., hand movement) are highlighted. Through image processing operations like dilation and region labeling, bounding boxes are fitted to these changed areas. The model is then applied solely to the image regions within these bounding boxes, and the results are then overlaid onto the previous segmentation map. Since typically only a small portion of the image changes, this approach saves significant computation. Although this method may result in a minor accuracy drop due to the lack of global context when processing local regions, it is not a primary concern here. We focus on presenting the method, with the possibility that future segmentation models could incorporate global context even in localized regions. Our preliminary, unoptimized results show an average of 2.44 FPS using this approach, compared to 1.69 FPS when the full frame is processed by the model. For more details, please refer to the accompanying code.
Figure 7: The approach for the second proposed solution is illustrated as follows: first, a binary change map is generated to highlight only the altered pixels, a computationally inexpensive process. Next, bounding boxes are fitted to these changed regions, and the model is applied exclusively to those areas. The segmentation map is shown here for illustrative purposes only, as the model used has not been specifically trained for human faces and bodies, which explains the lack of accuracy. The change map is computed similar to the first approach shown in Figure 2.
@@ 5 Conclusion
We highlighted a key issue with existing deep learning approaches and proposed solutions that can be integrated into current models or used to design new models with inherent attention and memory mechanisms. This work serves as a proof of concept and can be applied to other problems such as object detection, scene segmentation, and action recognition. It also has the potential to help address adversarial examples [2]. This method is particularly effective when the input has high resolution.
@@ 6 Appendix
Convolution and pooling modules of CNNs are shown in Figures 89, and  10.